### v1

# Model-free Rayleigh weight from x-ray Thomson scattering measurements

Abstract
X-ray Thomson scattering (XRTS) has emerged as a powerful tool for the diagnostics of matter under extreme conditions. In principle, it gives one access to important system parameters such as the temperature, density, and ionization state, but the interpretation of the measured XRTS intensity usually relies on theoretical models and approximations. In this work, we show that it is possible to extract the Rayleigh weight—a key property that describes the electronic localization around the ions—directly from the experimental data without the need for any model calculations or simulations. As a practical application, we consider an experimental measurement of strongly compressed Be at the National Ignition Facility (NIF) [Döppner et al., Nature 618, 270-275 (2023)]. In addition to being interesting in their own right, our results will open up new avenues for diagnostics from ab initio simulations, help to further constrain existing chemical models, and constitute a rigorous benchmark for theory and simulations.
The x-ray Thomson scattering (XRTS) method Glenzer and Redmer (2009); Sheffield et al. (2010) constitutes a powerful experimental technique, which is capable of giving microscopic insights into a probed sample. A particularly important use case for XRTS is the diagnostics of experiments with matter under extreme densities, temperatures and pressures Falk (2018). Such warm dense matter (WDM) Graziani et al. (2014); Dornheim et al. (2023a) naturally occurs in various astrophysical objects such as giant planet interiors Benuzzi-Mounaix et al. (2014); Guillot et al. (2022); Drake (2018), brown dwarfs Becker et al. (2014) and the outer crust of neutron stars Gudmundsson et al. (1983). Moreover, WDM plays an important role in a variety of technological applications such as inertial confinement fusion (ICF) Betti and Hurricane (2016), where the fuel capsule has to traverse the WDM regime in a controlled way to reach ignition Hu et al. (2011). The recent breakthrough at the National Ignition Facility (NIF) to reach ignition The Indirect Drive ICF Collaboration (2024) has further substantiated the importance of accurately diagnosing WDM states. In the laboratory, these extreme states can be realized using a variety of experimental techniques Falk (2018), and XRTS is often used to infer a-priori unknown system parameters such as the mass density $\rho$, temperature $T$, and ionization state $Z$; see, e.g., Ref. Plagemann et al. (2015). These properties can then be used for physical considerations, to inform equation-of-state tables Falk et al. (2012, 2014); Gaffney et al. (2018), and to benchmark integrated multi-scale simulations such as radiation hydrodynamics Pomraning (2005).
In practice, the measured XRTS intensity can be accurately expressed as Gawne et al. (2024a)
$I(\mathbf{q},E)=S_{ee}(\mathbf{q},E)\circledast R(E)\ ,$ (1)
where $S_{ee}(\mathbf{q},E)$ denotes the electronic dynamic structure factor (DSF) that describes the probed system, and $R(E)$ the combined source-and-instrument function (SIF) that takes into account the shape of the x-ray source and effects of the detector Gawne et al. (2024a). We note that a deconvolution solving Eq. (1) for $S_{ee}(\mathbf{q},E)$ constitutes a major challenge considering the uncertainties of the SIF and the experimental error bars Höll et al. (2007); Fortmann et al. (2009). Therefore, the usual way to interpret the measured XRTS intensity has been to construct a forward model for $S_{ee}(\mathbf{q},E)$, convolve it with $R(E)$, and then fit the resulting trial intensity to the LHS of Eq. (1) where the unknown system properties are treated as free parameters. Naturally, the thus inferred information depends on the employed forward model for $S_{ee}(\mathbf{q},E)$. These are usually based on a number of assumptions such as the possibility to distinguish between bound and free electrons as proposed within the popular Chihara approach Chihara (1987); Höll et al. (2007); Fortmann et al. (2009); Glenzer and Redmer (2009); Kraus et al. (2019); Döppner et al. (2023); Böhme et al. (2023).
Figure 1: Left: XRTS measurement on strongly compressed Be at the NIF Döppner et al. (2023) (green) and its decomposition into elastic (grey) and inelastic (red) contributions. Center: Determination of the normalization $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$ from the f-sum rule Dornheim et al. (2024a), cf. Eq. (7). Right: Wavenumber dependence of the Rayleigh weight $W_{R}(\mathbf{q})$, comparing the experimental data point [Eq. (6)] with ab initio PIMC simulations Dornheim et al. (2024b); the solid lines have been obtained for $T=155.5\,$eV and the shaded areas correspond to the uncertainty range of ±$\pm 15\,$eV.
Very recently, it has been proposed to instead consider the two-sided Laplace transform of the DSF Dornheim et al. (2022, 2023b, 2023a),
$F_{ee}(\mathbf{q},\tau)=\mathcal{L}\left[S_{ee}(\mathbf{q},E)\right]=\int_{-\infty}^{\infty}\textnormal{d}E\ S_{ee}(\mathbf{q},E)\ e^{-\tau E}\ ,$ (2)
which is directly related to the imaginary-time density–density correlation function (ITCF) $F_{ee}(\mathbf{q},\tau)$. The latter naturally emerges in Feynman’s path-integral picture of statistical mechanics Kleinert (2009) and, from a physical perspective, contains the same information as $S_{ee}(\mathbf{q},E)$, although in an unfamiliar representation Dornheim et al. (2023c, d). It corresponds to the usual intermediate scattering function $F(\mathbf{q},t)=\braket{\hat{n}(\mathbf{q},t)\hat{n}(-\mathbf{q},0)}$ with an imaginary argument $t=-i\hbar\tau$ with $\tau\in[0,\beta]$ and $\beta=1/k_{\textnormal{B}}T$ the inverse temperature. A key advantage of working in the Laplace domain is given by the convolution theorem
$\mathcal{L}\left[S_{ee}(\mathbf{q},E)\right]=\frac{\mathcal{L}\left[S_{ee}(\mathbf{q},E)\circledast R(E)\right]}{\mathcal{L}\left[R(E)\right]}\ ,$ (3)
which is remarkably stable with respect to experimental noise Dornheim et al. (2023b, 2024a). Eq. (3) thus gives one direct access to $F_{ee}(\mathbf{q},\tau)$ from the experimental observation. As a first application, Dornheim et al. Dornheim et al. (2022) have suggested to consider the imaginary-time version of the detailed balance relation $F_{ee}(\mathbf{q},\tau)=F_{ee}(\mathbf{q},\beta-\tau)$, which gives one model-free access to the temperature of arbitrarily complex systems Dornheim et al. (2022, 2023b); Böhme et al. (2023); Schörner et al. (2023). A second practical application of the ITCF in the context of XRTS diagnostics is given by the f-sum rule, which relates the first $\tau$-derivative of $F_{ee}(\mathbf{q},\tau)$ to the momentum transfer $q=|\mathbf{q}|$ that follows from the scattering angle $\theta$ Dornheim et al. (2024a, 2023e). In this way, one can infer the normalization of the measured intensity, which is given by the electronic static structure factor $S_{ee}(\mathbf{q})$—the Fourier transform of the electron–electron pair correlation function $g_{ee}(\mathbf{r})$. Finally, we mention the recent idea by Vorberger et al. Vorberger et al. (2024), who have proposed to utilize Eq. (3) to quantify the degree of electronic non-equilibrium in the probed system.
In the present work, we extend these efforts towards a model-free diagnostics of XRTS measurements by extracting the Rayleigh weight $W_{R}(\mathbf{q})=S_{eI}^{2}(\mathbf{q})/S_{II}(\mathbf{q})$ [where $S_{eI}(\mathbf{q})$ and $S_{II}(\mathbf{q})$ are the electron–ion and ion–ion static structure factor], which describes the degree of electronic localization around the ions Vorberger and Gericke (2015); Ma et al. (2013), directly from the experimental observation. Our idea is based on a combination of the f-sum rule with the ratio of elastic to inelastic scattering $r(\mathbf{q})$, cf. Eq. (5) below. Therefore, it is generally available and even extends to non-equilibrium situations. As a practical example, we consider an XRTS measurement on strongly compressed Be that has been carried out at the NIF Döppner et al. (2023). The proposed direct inference of $W_{R}(\mathbf{q})$ helps to further constrain forward models for the interpretation of XRTS measurements, which is very important in its own right. In addition, we expect $W_{R}(\mathbf{q})$ to be a valuable diagnostic tool. For example, one might first infer the temperature $T$ from a given XRTS signal using the model-free ITCF method Dornheim et al. (2022) and then carry out a set of ab intio calculations for $W_{R}(\mathbf{q})$ over a reasonable interval of densities $\rho$; here, we use highly accurate path integral Monte Carlo (PIMC) simulations using the set-up described in Ref. Dornheim et al. (2024b) and density functional theory molecular dynamics (DFT-MD) simulations by Bethkenhagen et al. Bethkenhagen et al. . We find excellent agreement between both methods, whereas computationally less involved average-atom models Sterne et al. (2007) and Chihara fits deviate. This is of direct practical consequence for the interpretation of the experimental signal and suggests a significantly lower density of $\rho=(22\pm 2)$ g/cm3 (see also Ref. Dornheim et al. (2024b)) compared to the Chihara based estimate of $\rho=(34\pm 4)$ g/cm3 by Döppner et al. Döppner et al. (2023).
From a methodological perspective, we note that $W_{R}(\mathbf{q})$ constitutes a perfect observable for DFT-MD, as it only involves the static single-electron density distribution $n_{e}(\mathbf{r})$ [and the ion density distribution and correlation function, which are both easily accessible within MD]. This is in contrast to time-dependent DFT (TD-DFT) calculations that, in practice, contain additional approximations such as the unknown dynamic exchange–correlation kernel in the case of linear-response TD-DFT Ullrich (2012); Moldabekov et al. (2023a, b); Schörner et al. (2023). Finally, we mention the possibility to use experimental results for $W_{R}(\mathbf{q})$ as a rigorous benchmark for simulations and theoretical models in situations where the density and temperature are already being known by other means.
Idea. In general, we can split the DSF into a quasi-elastic contribution $S_{\textnormal{el}}(\mathbf{q},E)$ that is due to the electronic localization around the ions described by the Rayleigh weight, and the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ [cf. Fig. 1a)],
$S_{ee}(\mathbf{q},E)$ $=$ $\underbrace{W_{R}(\mathbf{q})\delta(E)}_{S_{\textnormal{el}}(\mathbf{q},E)}+S_{\textnormal{inel}}(\mathbf{q},E)\ .$ (4)
Here, the quasi-elastic nature of the first part is due to the substantially longer ionic time-scales. In practice, treating $S_{\textnormal{el}}(\mathbf{q},E)$ as a delta distribution is appropriate if the SIF $R(E)$ is significantly broader than the actual ion feature. This is usually the case except for dedicated experiments that aim to resolve ionic energy scales on purpose Descamps et al. (2020). Within the chemical picture that assumes a decomposition into effectively bound and free electrons Glenzer and Redmer (2009); Gregori et al. (2003); Kraus et al. (2019); Döppner et al. (2023), the inelastic part $S_{\textnormal{inel}}(\mathbf{q},E)$ consists of both bound-free transitions (and their reverse process Böhme et al. (2023)) and free-free transitions. The Rayleigh weight thus constitutes an indispensable ingredient to Chihara models Chihara (1987); Gregori et al. (2003) that are widely used for the interpretation of XRTS experiments with WDM Böhme et al. (2023); Döppner et al. (2023); Kraus et al. (2019); Glenzer and Redmer (2009). However, it is important to note that we do not have to make such an approximate distinction between bound and free electrons in the present work.
Let us next consider the ratio of the elastic and inelastic contributions
$r(\mathbf{q})=\frac{\int_{-\infty}^{\infty}\textnormal{d}\omega\ S_{\textnormal{el}}(\mathbf{q},\omega)}{\int_{-\infty}^{\infty}\textnormal{d}\omega\ S_{\textnormal{inel}}(\mathbf{q},\omega)}=\frac{W_{R}(\mathbf{q})}{S_{ee}(\mathbf{q})-W_{R}(\mathbf{q})}\ ,$ (5)
that constitutes a standard observable in XRTS experiments Döppner et al. (2023), see, e.g., Fig. 1a). Here $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$ denotes the aforementioned electronic static structure factor that we can also directly extract from the XRTS signal using the f-sum rule applied in the imaginary-time domain as it has been explained in detail in the recent Ref. Dornheim et al. (2024a), see also Eq. (7) below. Solving Eq. (5) for the Rayleigh weight then gives
$W_{R}(\mathbf{q})=\frac{S_{ee}(\mathbf{q})}{1+r^{-1}(\mathbf{q})}\ .$ (6)
Figure 2: Comparison of simulation results for the Rayleigh weight at $\rho=20\,$g/cc [top] and $\rho=30\,$g/cc [bottom]. Blue cross: experiment Döppner et al. (2023); solid green: PIMC; dashed red: average atom model; black and yellow dotted: chemical model with ionization states of $Z=3.52$ and $Z=3.73$; dash-dotted purple: DFT-MD Bethkenhagen et al. . PIMC and CM have been computed for $T=155.5\,$eV, DFT-MD and AA for $T=150\,$eV.
Figure 3: Curves: DFT-MD results for $W_{R}(\mathbf{q})$ at $T=150\,$eV for different mass densities $\rho$. The experimental measurement (blue cross) at the NIF Döppner et al. (2023) is associated with a density of $\rho=22\pm 2\,$g/cc [dash-dotted pink], see also the inset showing a magnified segment.
Results. In Fig. 1, we apply our new idea to an XRTS data set that has been recently obtained at the NIF by Döppner et al. Döppner et al. (2023). Panel a) shows the full XRTS intensity [solid green, cf. Eq. (1)], which can easily be decomposed into its elastic (grey) and inelastic (red) contribution if the SIF $R(E)$ is known; this is indeed the case for the backlighter x-ray source utilized at the NIF MacDonald et al. (2021). The second ingredient to Eq. (6) is given by the electronic static structure factor $S_{ee}(\mathbf{q})$, which we obtain from the Laplace transform of the XRTS signal via the f-sum rule. This procedure is illustrated in Fig. 1b), where we show the deconvolved $F_{ee}(\mathbf{q},\tau)$ as a function of $\tau$. In the imaginary-time domain, the f-sum rule is given by Dornheim et al. (2023e, c, 2024a)
$\left.\frac{\partial}{\partial\tau}F_{ee}(\mathbf{q},\tau)\right|_{\tau=0}=-\frac{\hbar^{2}\mathbf{q}^{2}}{2m_{e}}\ ;$ (7)
matching Eq. (7) [dashed red line] with the LHS. of Eq. (3) [solid blue line] around $\tau=0$ then determines the a-priori unknown normalization constant and, in turn, $S_{ee}(\mathbf{q})=F_{ee}(\mathbf{q},0)$. Inserting $r(\mathbf{q})$ and $S_{ee}(\mathbf{q})$ into Eq. (6) gives our final, model-free estimate of the Rayleigh weight, which is shown as the blue cross in Fig. 1c) at the corresponding experimental wavenumber of $q=7.89\,$Å-1. The red, green and yellow solid curves in the panel show highly accurate ab initio PIMC results that have been obtained for $N_{\textnormal{Be}}=10$ Be atoms using the set-up described in the recent Refs. Dornheim et al. (2024b, c) at the temperature of $T=155.5$ eV that has been inferred from the symmetry of the ITCF; see the appendix of Ref. Dornheim et al. (2024b) for additional details. In addition, the associated colored areas correspond to PIMC simulations for $T=(155.5\pm 15)$ eV and indicate the uncertainty in the Rayleigh weight at a given density due to the uncertainty in the inferred temperature. Fig. 1 thus indicates a mass density of $\rho\approx 20$ g/cm3, whereas the nominal result of $\rho=(34\pm 4)$ g/cm3 reported in the original Ref. Döppner et al. (2023) is ruled out.
To get further insights into the sensitivity of the interpretation of a given XRTS measurement to the employed model or simulation technique, we compare a number of independent methods in Fig. 2, where the top and bottom panels correspond to $\rho=20$ g/cm3 and $\rho=30$ g/cm3, respectively. For both cases, we find very good agreement between PIMC (solid green) and DFT-MD simulations by Bethkenhagen et al. Bethkenhagen et al. (dash-dotted purple). The computationally cheaper average-atom model Sterne et al. (2007) (dashed red) is in qualitative, though not quantitative agreement with the ab initio data sets. Instead, it agrees with the experimental data point for $\rho=30$ g/cm3. This clearly highlights the sensitivity of equation-of-state measurements to the proper treatment of electronic correlations and other many-body effects in the analysis of the experimental observation. Finally, we include two chemical models for the hypothetical ionization degrees of $Z=3.52$ (dotted black) and $Z=3.73$ (dotted yellow) for both densities, see the Supplemental Material for additional details sup . Evidently, the inferred density strongly depends on the inferred ionization state, making the invocation of additional constraints desirable. In Ref. Döppner et al. (2023), Döppner et al. have used the Saha equation in combination with a semi-empirical form factor lowering model for this purpose, leading to the nominal parameters of $\rho=(34\pm 4)$ g/cm3 and $Z=3.4\pm 0.1$. A systematic investigation of the origin of this discrepancy between their chemical model and the ab initio data sets from PIMC and DFT-MD is beyond the scope of the present work and will be pursued in a dedicated future study.
Let us conclude with a systematic analysis of the dependence of $W_{R}(\mathbf{q})$ on the density that is shown in Fig. 3 based on extensive DFT-MD results for $\rho=(5-80)$ g/cm3 at $T=150$ eV. The main trend is given by the systematic increase of the Rayleigh weight with $\rho$, which is to a large degree due to the different length scales in the system; the effect almost vanishes when one adjusts the $x$-axis by the Fermi wavenumber $q_{\textnormal{F}}\sim 1/\rho^{1/3}$. The inset shows a magnified segment around the experimental data point, with the latter being located between the DFT-MD results for $\rho=20$ g/cm3 (red) and $\rho=24$ g/cm3 (black); this leads to our final estimate for the density of $\rho=(22\pm 2)$ g/cm3 [dash-dotted pink].
Summary and Discussion. We have presented a new approach for the model-free extraction of the Rayleigh weight $W_{R}(\mathbf{q})$ from XRTS measurements. Most importantly, $W_{R}(\mathbf{q})$ constitutes an important measure for the electronic localization around the ions, which is interesting in its own right. In addition, $W_{R}(\mathbf{q})$ is of direct practical value for XRTS diagnostics, as we have demonstrated in detail on a recent experiment with strongly compressed beryllium at the NIF Döppner et al. (2023). In particular, we propose to first infer the temperature $T$ from the model-free ITCF thermometry approach Dornheim et al. (2022, 2023b), and to subsequently match the experimental result for $W_{R}(\mathbf{q})$ with simulation results over a reasonable range of densities. We note that $W_{R}(\mathbf{q})$ is a particularly suitable observable for DFT-MD simulations, as it does not involve any dynamic information such as the a-priori unknown dynamic exchange–correlation kernel. We find excellent agreement between DFT-MD and PIMC reference data; this is encouraging since direct PIMC simulations are still limited to low-$Z$ materials at moderate to high temperatures Dornheim (2019); Bonitz et al. (2024), whereas DFT-MD is more broadly applicable. An additional advantage of $W_{R}(\mathbf{q})$ over alternative observables such as the ratio of elastic to inelastic scattering $r(\mathbf{q})$ is that the former does not require any explicit information about the electronic static structure factor $S_{ee}(\mathbf{q})$, which is notoriously difficult for DFT-based methodologies.
In practice, we infer a mass density of $\rho=(22\pm 2)$ g/cm3 from the beryllium data set using either PIMC or DFT-MD simulations, which is significantly lower than the nominal value of $\rho=(34\pm 4)$ g/cm3 that has been reported in the original publication based on a chemical Chihara model. Moreover, we find that the computationally cheaper average-atom model tends to overestimate the density, whereas chemical models generally require additional constraints for parameters such as the ionization degree. Our study thus clearly highlights the importance of an accurate treatment of quantum many-body effects for XRTS based equation-of-state measurements even at relatively high temperatures; the main value of average-atom calculations is their low computational efforts, making them a useful tool to constrain the parameter space for more expensive ab initio simulations.
We are convinced that our work opens up new possibilities for the study of warm dense matter and beyond. An important point for future research is given by the sensitivity of the inferred Rayleigh weight to the SIF $R(E)$, which is usually modelled for backlighter set-ups MacDonald et al. (2021), but can be known with high precision at modern XFEL facilities Gawne et al. (2024a). Indeed, recent advances in high-resolution XRTS measurements Gawne et al. (2024b) at the European XFEL will likely facilitate the application of the model-free ITCF thermometry technique even at moderate temperatures of $T\sim 1\,$eV, which are of relevance for both planetary and material science; the current model-free framework for the extraction of $W_{R}(\mathbf{q})$ is applicable at any temperature. In combination, these two methods will allow for accurate equation-of-state measurements in previously inaccessible regimes. Finally, we mention the intriguing possibility of performing an XRTS experiment on an isochorically heated sample with an appropriate delay between pump and probe to ensure proper equilibration. Since both $T$ and $\rho$ would be well known in such a scenario, a corresponding measurement of $W_{R}(\mathbf{q})$ using the present framework would constitute a truly unambiguous reference data set for the rigorous benchmarking of ab initio simulations and chemical models alike.
Acknowledgments
This work was partially supported by the Center for Advanced Systems Understanding (CASUS), financed by Germany’s Federal Ministry of Education and Research (BMBF) and the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Union’s Just Transition Fund (JTF) within the project Röntgenlaser-Optimierung der Laserfusion (ROLF), contract number 5086999001, co-financed by the Saxon state government out of the State budget approved by the Saxon State Parliament. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2022 research and innovation programme (Grant agreement No. 101076233, ”PREXTREME”). This work was partially supported by the German Research Foundation (DFG) within the Research Unit FOR 2440. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Computations were performed on a Bull Cluster at the Center for Information Services and High-Performance Computing (ZIH) at Technische Universität Dresden and at the Norddeutscher Verbund für Hoch- und Höchstleistungsrechnen (HLRN) under grants mvp00018 and mvp00024. SH was supported by Sandia National Laboratories, a multi-mission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for DOE’s National Nuclear Security Administration under contract DE-NA0003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.